It is known that heat is added to a gas in sealed container. The container is fitted with a moveable piston. It is known that heat is added to a gas in sealed container. The container is fitted with a moveable piston.

The temperature of the gas decreases.

The temperature of the gas increases.

The temperature of the gas remains the same.

Not enough information is given to answer this question.

(Temp of gas increases was incorrect)
Answer #1

If the piston remains stationary then the work done by the gas will be zero , W = 0

Then U = Q - W

cv (dT) = Q ( since heat is added, temperature increases)

If the piston moves and the  the work done by the gas is greater than the heat added, then

U = Q - W will be less than 0

Hence the temperature will decrease

If the piston moves and the  the work done by the gas is equal to the heat added, then

U = Q - W = 0

Hence the temperature remains same

Since it is not mentioned in the question about work done by the gas on moveable piston, hence Not enough information is given to answer this question.